# Anchored Preference Optimization and Contrastive Revisions: Addressing Underspecification in Alignment

### v2

@@ 1 Introduction
In order to study the role of both (i) minimally contrastive preference data, and (ii) distinct alignment training dynamics, we individually align a model across four comparable preference datasets using five alignment objectives. One dataset is created through our CLAIR method. We compare this with two conventional judge-based datasets (Reinforcement Learning from AI Feedback; Bai et al. 2022). Finally, we consider an ablated version of CLAIR created to directly assess the impact of contrastiveness. We consider five distinct alignment objectives: DPO (Rafailov et al., 2024b), KTO (Ethayarajh et al., 2024), continued Supervised Fine-Tuning on the preferred answer, and two variants of our proposed APO. We measure MixEval-Hard accuracy (Ni et al., 2024) and length-controlled AlpacaEval scores (Dubois et al., 2024) for each model, both correlate highly with model rankings produced by humans (Chiang et al., 2024).
@@ 3 Contrastive Learning from Revisions
Table 1: Average token-level Jaccard similarity (intersection over union) and average character-level Levenshtein edit-distance between winning $y_{w}$ and losing $y_{l}$ answers for four comparable preference datasets built on top of Llama-3-8B-Instruct. The CLAIR dataset is the most contrastive on both metrics.
@@ 5.1 Evaluation Methodology
At the time of writing, MixEval-Hard (Ni et al., 2024) is the benchmark with the highest Chatbot Arena correlation (0.98 rank correlation). MixEval-Hard features hard queries with known answers across a wide range of domains and uses a GPT3.5-turbo (Brown et al., 2020; Ouyang et al., 2022) model to evaluate if predicted answers correspond with this ground-truth. This makes MixEval-Hard more grounded in human knowledge and significantly cheaper to run compared to other popular evaluation frameworks such as AlpacaEval (Li et al., 2023; Dubois et al., 2024). Under the hood, MixEval-Hard utilizes queries sampled from MATH (Hendrycks et al., 2021), BBH (Suzgun et al., 2023), DROP (Dua et al., 2019), GSM8k (Cobbe et al., 2021), AGIEval (Zhong et al., 2024), TriviaQA (Joshi et al., 2017), MBPP (Austin et al., 2021), MMLU, (Hendrycks et al., 2020), HellaSwag (Zellers et al., 2019), BoolQ (Clark et al., 2019), GPQA (Rein et al., 2023), PIQA (Bisk et al., 2020), OpenBookQA (Mihaylov et al., 2018), ARC (Clark et al., 2018), CommonsenseQA (Talmor et al., 2019), and SIQA (Sap et al., 2019).
@@
Additionally, we consider the length-controlled LC-AlpacaEval2.0 win rate (Dubois et al., 2024) as an auxiliary evaluation method. LC-AlpacaEval2.0 measures how often a generated output is preferred to a reference output, as judged by a GPT4 model across a wide set of instructions. On this benchmark, Llama-3-8B-Instruct attains a win rate of 22.92% whereas GPT4 has a win rate of 50%.
@@
Table 2: Max and mean MixEval-Hard improvements for 2024-06-01 and 2024-08-11 splits, aggregated over 18 epochs of aligning Llama-3-8B-Instruct. Best overall performance bold, best performance per dataset underlined, standard deviation in parentheses. While MixEval-Hard correlates better with human model ratings, we also report the average LC-AlpacaEval2.0 score increase over the two best MixEval-Hard checkpoints, and average length increase (in characters) of the responses. CLAIR leads to the greatest overall performance improvement on MixEval-Hard. APO methods achieve the best performance across both Judged and CLAIR datasets.
@@ 5.3 Results
We report the maximal and mean MixEval-Hard improvement over all checkpoints from the same training run. This helps us understand both the best-case and average impact of alignment across the entire training procedure. We use both 2024-06-01 and 2024-08-11 versions of MixEval-Hard, which each feature a distinct set of queries. Due to the increased cost associated with LC-AlpacaEval2.0, we only measure the win rate for the two best MixEval-Hard checkpoints and report their average. This analysis is summarized in Table 2 for every dataset and objective; we now discuss these results in more detail.
@@ 5.3.1 Preference Data
We noted in Section 1 that uncontrolled contrastiveness can degrade model performance. We see this dramatically in the results for the Stronger Preferred dataset, which heavily degrades model performance. Like CLAIR, this dataset has all winning outputs produced by a stronger model. Unlike CLAIR, though, its examples provide no guarantee of relevant minimal contrasts. Thus, the contrastiveness induced by the CLAIR revision process is a major driver of performance.
@@
For the LC-AlpacaEval2.0 benchmark, on-policy judge leads to better performance than CLAIR. However, LC-AlpacaEval2.0 correlates less with human-produced model rankings than MixEval-Hard.
A breakdown of performance in function of MixEval-Hard’s constituent benchmarks is given in Appendix B for every dataset. We remark on a potential issue we have encountered with LC-AlpacaEval2.0 in Appendix C.
@@ 5.3.2 Alignment Objectives
On MixEval-Hard, Anchored Preference Optimization (APO) consistently leads to the greatest performance increase for every preference dataset, with the exception of the Stronger Preferred dataset, where all contrastive objectives deteriorate the model. The relation between the preference dataset and the target model controls which variant of APO is best for any dataset, as predicted in Section 2. APO-down results in the best performance when winning outputs are generally worse than the target model, as is the case for the off-policy judge dataset. APO-zero is the best objective when winning outputs are generally better than the target model, as is the case for CLAIR and on-policy judge datasets. The difference between alignment objectives is less salient for the on-policy judge dataset as compared to CLAIR, since winning on-policy judge outputs are only slightly better than Llama-3-8B-Instruct on average. Winning CLAIR outputs may be vastly better than Llama-3-8B-Instruct since they are produced by a stronger model, making the different in alignment objectives more noticeable.
@@ 6 Related Work
We articulated two core insights concerning (i) the role of contrastive preference data, and (ii) the need to anchor alignment depending on model and data. These insights translate to any effort which uses comparative preferences.
For the remainder of this review, we focus on contrastive methods which directly align an LM on preference data (of which Wang et al. 2024 provide a detailed overview).
@@
Ethayarajh et al. (2024) and Richemond et al. (2024) use unpaired examples and rewards for alignment instead of paired examples. Zhang et al. (2024) and Duan et al. (2024) operate solely on undesirable examples in this unpaired setting. In contrast, our work exclusively operates on paired preferences. However, the core insights of APO do apply to unpaired data. For example, Ethayarajh et al. (2024) use binary desired / undesired labels for each answer. We argue this desirability is inherently relative to the model: the same example of desirable behavior used to improve a weak model may actually be an example of undesirable behavior compared to a stronger model, causing the need for anchoring.
@@ 7 Conclusion
Alignment performance is significantly impacted by (i) the contrastiveness of the preference pairs and (ii) the relationship between model and data . We introduce Contrastive Learning from AI Revisions (CLAIR), a data-creation method which produces more contrastive preference pairs, and Anchored Preference Optimization (APO), a family of alignment objectives with tailored training dynamics. Our experiments aligning Llama-3-8B-Instruct show that CLAIR preferences lead to the highest performance improvement, and APO methods consistently outperform conventional alignment objectives.
@@ Appendix C How well does AlpacaEval control for response lengths?
GPT4 is known to favor more verbose responses, which can artificially inflate AlpacaEval win rates for verbose models (Dubois et al., 2024). To counteract this bias, Dubois et al. (2024) estimate a length-controlled AlpacaEval win rate, which we report on in Table 2. Specifically, the authors adopt a causal inference framework to answer the question "What would the AlpacaEval metric be, if the outputs of all models had the same length as those of the baseline?" (Dubois et al., 2024).
In order to meaningfully apply causal inference, a few key assumptions need to be met. The Positivity assumption (Hernán and Robins, 2006) states that, when estimating the effect of a treatment (e.g., the effect of response length on a model’s win rate), for all covariates (e.g., for all models), there are at least some subjects which receive the treatment (e.g., there are some responses of that model which are long and some which are not). Intuitively, the Positivity assumption applied to the length-control question states that you need to observe at least some long and some short responses for every model in order to accurately estimate how the response length influences the model’s win rate.
@@
One potential avenue towards a more robust length-controlled win rate would be to specifically prompt models to generate shorter or longer answers if the Positivity constraint is not met. Unfortunately, a more thorough study of length-controlled win rate is out of scope for this work.

### v3

@@ 1 Introduction
In order to study the role of both (i) minimally contrastive preference data, and (ii) distinct alignment training dynamics, we individually align a model across four comparable preference datasets using five alignment objectives. One dataset is created through our CLAIR method. We compare this with two conventional judge-based datasets (Reinforcement Learning from AI Feedback; Bai et al. 2022). Finally, we consider an ablated version of CLAIR created to directly assess the impact of contrastiveness. We consider five distinct alignment objectives: DPO (Rafailov et al., 2024b), KTO (Ethayarajh et al., 2024), continued Supervised Fine-Tuning on the preferred answer, and two variants of our proposed APO. We measure MixEval-Hard accuracy (Ni et al., 2024) and length-controlled AlpacaEval scores (Dubois et al., 2024) for each model, both benchmarks correlate highly with model rankings produced by humans (Chiang et al., 2024).
@@ 3 Contrastive Learning from Revisions
Table 1: Average token-level Jaccard similarity (intersection over union) and average character-level Levenshtein edit-distance between winning $y_{w}$ and losing $y_{l}$ answers for four comparable preference datasets built on top of Llama-3-8B-Instruct. The CLAIR dataset produces the best contrasts on both metrics.
@@ 5.1 Evaluation Methodology
MixEval-Hard (Ni et al., 2024) is a benchmark with very high Chatbot Arena correlation (0.96 rank correlation). MixEval-Hard features hard queries with known answers across a wide range of domains and uses a GPT3.5-turbo (Brown et al., 2020; Ouyang et al., 2022) model to evaluate if predicted answers correspond with this ground-truth. This makes MixEval-Hard more grounded in human knowledge and significantly cheaper to run compared to other popular evaluation frameworks such as AlpacaEval (Li et al., 2023; Dubois et al., 2024). Under the hood, MixEval-Hard utilizes queries sampled from MATH (Hendrycks et al., 2021), BBH (Suzgun et al., 2023), DROP (Dua et al., 2019), GSM8k (Cobbe et al., 2021), AGIEval (Zhong et al., 2024), TriviaQA (Joshi et al., 2017), MBPP (Austin et al., 2021), MMLU, (Hendrycks et al., 2020), HellaSwag (Zellers et al., 2019), BoolQ (Clark et al., 2019), GPQA (Rein et al., 2023), PIQA (Bisk et al., 2020), OpenBookQA (Mihaylov et al., 2018), ARC (Clark et al., 2018), CommonsenseQA (Talmor et al., 2019), and SIQA (Sap et al., 2019).
@@
Additionally, we consider the length-controlled LC-AlpacaEval2.0 win rate (Dubois et al., 2024). However, two factors lead us to favor MixEval-Hard as our primary evaluation tool. The first is practical: LC-AlpacaEval2.0 is prohibitively expensive to run, we thus use MixEval-Hard for the bulk of our evaluation. The second concerns the assessment itself: while both benchmarks are highly correlated with human-produced model rankings, MixEval-Hard utilizes questions with known ground-truth answers whereas LC-AlpacaEval2.0 uses an LLM judge without any ground-truth to decide correctness.
@@
Table 2: Max and mean MixEval-Hard improvements for the 2024-06-01 and 2024-08-11 splits, aggregated over 18 epochs of aligning Llama-3-8B-Instruct. Best overall performance bold, best performance per dataset underlined, standard deviation in parentheses. While MixEval-Hard functions as our primary evaluation tool, we also report the average LC-AlpacaEval2.0 score increase over the two best MixEval-Hard checkpoints, and average length increase (in characters) of the responses. CLAIR leads to the greatest overall performance improvement on MixEval-Hard. APO methods achieve the best performance across both Judged and CLAIR datasets.
@@ 5.3 Results
We report the maximal and mean MixEval-Hard improvement over all checkpoints from the same training run. This helps us understand both the best-case and average impact of alignment across the entire training procedure. We use both 2024-06-01 and 2024-08-11 versions of MixEval-Hard, which each feature a distinct set of queries. Due to the increased cost associated with LC-AlpacaEval2.0, we only measure the win rate for the two best MixEval-Hard checkpoints and report their average. We use no system prompt for both evaluations. Our analysis is summarized in Table 2 for every dataset and objective; we now discuss these results in more detail.
@@ 5.3.1 Preference Data
We noted in Section 1 that uncontrolled contrastiveness can degrade model performance. We see this dramatically in the results for the Stronger Preferred dataset, which can heavily degrade model performance. Like CLAIR, this dataset has all winning outputs produced by a stronger model. Unlike CLAIR, though, its examples provide no guarantee of relevant minimal contrasts. Thus, the contrastiveness induced by the CLAIR revision process is a major driver of performance.
@@
The LC-AlpacaEval2.0 results generally follow a similar trend compared to MixEval-Hard, although the on-policy judge dataset attains a higher score compared to CLAIR. While both benchmarks correlate highly with human ratings of models, MixEval-Hard is our primary and most significant evaluation tool – we are able to evaluate every model checkpoint across two MixEval-Hard splits due to its low cost. Additionally, we remark on a potential issue with the robustness of LC-AlpacaEval2.0 in Appendix C. A performance breakdown in function of MixEval-Hard’s constituent benchmarks is given in Appendix B.
@@ 5.3.2 Alignment Objectives
On MixEval-Hard, Anchored Preference Optimization (APO) consistently leads to the greatest performance increase for every preference dataset, with the exception of the Stronger Preferred dataset, where all contrastive objectives underperform SFT. The relation between the preference dataset and the target model controls which variant of APO is best for any dataset, as predicted in Section 2. APO-down results in the best performance when winning outputs are generally worse than the target model, as is the case for the off-policy judge dataset. APO-zero is the best objective when winning outputs are generally better than the target model, as is the case for CLAIR and on-policy judge datasets. The difference between alignment objectives is less salient for the on-policy judge dataset as compared to CLAIR, since winning on-policy judge outputs are only slightly better than Llama-3-8B-Instruct on average. Winning CLAIR outputs may be vastly better than Llama-3-8B-Instruct since they are produced by a stronger model, making the different in alignment objectives more noticeable.
@@ 6 Related Work
We articulated two core insights concerning (i) the role of contrastive preference data, and (ii) the need to anchor alignment depending on model and data. These insights translate to any alignment effort which uses comparative preferences. For example, a reward model trained on spurious preference signals may be a less accurate proxy for real rewards, contributing to problems such as reward overoptimization or hacking (Gao et al., 2023; Rafailov et al., 2024a).
For the remainder of this review, we first focus on contrastive alignment methods and their variants (of which Wang et al. 2024 provide a detailed overview). Finally, we discuss related preference datasets and how they were created.
@@
Ethayarajh et al. (2024), Richemond et al. (2024), and Jung et al. (2024) use unpaired examples and rewards for alignment instead of paired examples. Zhang et al. (2024) and Duan et al. (2024) operate solely on undesirable examples in this unpaired setting. In contrast, our work exclusively operates on paired preferences. However, the core insights of APO do apply to unpaired data. For example, Ethayarajh et al. (2024) use binary desired / undesired labels for each answer. We argue this desirability is inherently relative to the model: the same example of desirable behavior used to improve a weak model may actually be an example of undesirable behavior compared to a stronger model, causing the need for anchoring.
@@
Preference Datasets:
Chiang et al. (2024) release a dataset of human preference judgements across conversations between humans and several AI assistants. To alleviate the need for human judges, some efforts focus on scaling preference annotations with LLM-based judges (Cui et al., 2024; Zhu et al., 2023) or metric-based judges (Jiang et al., 2023). Unlike our CLAIR method, these works do not create preferences through revisions. Bai et al. (2022) use a set of predetermined criteria (called a constitution) to prompt an LLM to revise answers and make them safer (see also Lambert et al. 2024). Dubey et al. (2024) used human revisions in the development of the llama-3.1 model family. While both efforts create preferences through revisions, we particularly focus on revisions that create a minimal contrast and studied the effect of this contrastiveness on alignment outcomes.
7 Future work
In this work, we have presented two variants of the APO objective family. Each method accounts for a distinct relationship between target model and preference pair during training. However, real world preference datasets may contain a wide range of different preference pairs, thus the dataset as a whole may not perfectly correspond with any single APO variant. To tackle this, a natural extension of APO could be to select the optimal APO variant at the preference pair level, instead of at the dataset level. Heuristically, this could be achieved using an off-the-shelf reward model to score each preference pair before training.
@@ 8 Conclusion
Alignment performance is significantly impacted by (i) the contrastiveness of the preference pairs and (ii) the relationship between target model and alignment data . We introduce Contrastive Learning from AI Revisions (CLAIR), a data-creation method which produces better contrasting preference pairs, and Anchored Preference Optimization (APO), a family of alignment objectives with tailored training dynamics. Our experiments aligning Llama-3-8B-Instruct show that CLAIR preferences lead to the highest performance improvement out of four comparable preference datasets, and APO methods consistently outperform conventional alignment objectives.
@@ Appendix C How well does AlpacaEval control for response lengths?
GPT4 as a judge is known to favor more verbose responses, which can artificially inflate AlpacaEval win rates for verbose models (Dubois et al., 2024). To counteract this bias, Dubois et al. (2024) estimate a length-controlled AlpacaEval win rate, which we report on in Table 2. Specifically, the authors adopt a causal inference framework to answer the question "What would the AlpacaEval metric be, if the outputs of all models had the same length as those of the baseline?" (Dubois et al., 2024).
In order to meaningfully apply causal inference, a few key assumptions need to be met. The Positivity assumption (Hernán and Robins, 2006) states that, when estimating the effect of a treatment, there are at least some subjects which receive the treatment for all covariates. Intuitively, the Positivity assumption applied to the length-control question states that you need to observe at least some long and some short responses for every model in order to accurately estimate how the response length influences the model’s win rate.
@@
While a more thorough study of length-controlled win rate is out of scope for this work, one potential avenue towards a more robust length-controlled win rate would be to specifically prompt models to generate shorter or longer answers if the Positivity constraint is not met.